Some conformally invariant fully nonlinear equations: Liouville, Yamabe and Harnack

Yanyan Li, Rutgers University

Abstract: I present some joint work with Aobing Li on some conformally invariant fully nonlinear equations. This includes works on Liouville type theorems, a fully nonlinear version of the Yamabe problem on locally conformally flat manifolds and some Harnack type inequalities.